David took his boat out for an all day ocean fishing trip. It took him five hours to reach his favorite fishing spot traveling with the current. The return trip to the harbor, however took him seven hours since he was going against the current. Assuming he set his cruise control to the same speed going out and coming back to the harbor and the speed of the ocean current that day was 5 miles per hour, at what speed was his boat set that day?

Answer by ewatrrr(24785)

D = rt
(r-5mph)7hr= (r+5mph)5hr
7r - 35 = 5r + 25
2r = 70
r = 35mph